Suppose the demand function facing a monopolist is given as p=50-q and the cost function is given as C=10q. Determine the total revenue function for the monopolist, and graph the relationship between total revenue, total cost and output. In your answer determine the profit maximizing level of output that would be chosen by the monopolist, and discuss mathematically how you know it is a profit maximum
Answer #1

Revenue - R - Demand x pnce Demand D P = 50-9 R =qxp R = q (502) IR = 504-92 Cost of product C = 109 The output is related to

Hope it helps you out.

In case any doubt please feel free to ask.

I will be happy to help you.

Thanks!!